OpenStudy (anonymous):

F(-5)=x^2+7x-35 can i get help please!!!!!!!!!!!!!!!!!!!!!!!

sam (.sam.):

F(x)=x^2+7x-35 F(-5)=(-5)^2+7(-5)-35 F(-5)=?

OpenStudy (anonymous):

what do it equal up to

sam (.sam.):

Compute the right hand side, F(-5)=(-5)^2+7(-5)-35 F(-5)=25-35-35 F(-5)=?